Prove that 1-cos2A/sin2A is tanA


1
Expert's answer
2021-02-24T06:02:52-0500

"\frac{1-cos2A}{sin2A}=\frac{1-(1-2sin^2A)}{2sinA*cosA}" ,


TRIGNOMETRIC DENTITY used : [Cos2A = 1-2sin2A]

[sin2A = 2sinA×cosA]


"\frac{1-cos2A}{sin2A}=\frac{1-1+2sin^2A}{2sinA\times cosA} = \frac{sinA}{cosA}=tanA"

So, "\frac{1-cos2A}{sin2A}=tanA"

LHS = RHS, Hence proved.